OpenStudy (anonymous):

Differentiate: y=5^(3x)

OpenStudy (anonymous):

Do you take the natural log of each side?

OpenStudy (blockcolder):

You know that \[\frac{d}{dx}a^x=a^x\ln{a}\]right? And you can rewrite \(5^{3x}\) as \((5^3)^x\).

OpenStudy (anonymous):

I got y'=5^(3x)*(ln5)

OpenStudy (blockcolder):

There's something missing there.

OpenStudy (anonymous):

Yeah its suppose to be 3ln5, but does that 3 come from when you take the derivative of 3x?

OpenStudy (blockcolder):

Yep, that's what you get when you apply the chain rule.

OpenStudy (anonymous):

thank you! How do you simplify dy/dx= y(2+2lnx)

OpenStudy (blockcolder):

What do you mean by simplify?

OpenStudy (anonymous):

The next step.. The problem is y=x^(2x)

OpenStudy (blockcolder):

You plug in y=x^(2x) into the equation for dy/dx.

OpenStudy (anonymous):

Seriously!? I've been here over thinking this. Thanks so much